Micheal is 52 years old today, making him 25 years older than Brain. Peter was 5 years younger than Brain and 10 years older than Jane five years ago. How old will jane me after 5 years from now?
Answer by mhaty(3)

52 - 25 = 27 ( Brain's age )
25 - 5 -5 = 15 ( Peter's age 5 years ago ) so his current age is 20
5 Years Ago :
Peter was 15 and Jane was 5.
so 5 + 5 = 10
adding 5 years from now, 10 + 5 = 15 :)